# Allocating Positional Goods: A Mechanism Design Approachthanks: I thank Krishna Dasaratha, Alex Gershkov, Bart Lipman, Chiara Margaria, and Juan Ortner for their helpful comments. The paper subsumes results in my note “A Mechanism Design Approach to ‘Gainers and Losers in Priority Services’.”

### v2

@@ 2.1 Setup
A monopolist seller (she) sells positional goods $q\in X\subseteq\mathbb{R}_{++}$ to a continuum of buyers (he) with unit mass. Denote by $q=0$ the outside option of not buying the good. Because consumers care about their relative position, consumption of positional goods induces buyers’ status (as defined later). Buyers are heterogeneous in their valuations of status (i.e., types), denoted by $\theta$. Assume the buyer’s type $\theta$ has a distribution $F(\theta)$ with full support $\Theta=[0,\bar{\theta}]$ and density $f(\theta)$.777 The support can be extended to $\bar{\theta}=\infty$ and $\underline{\theta}>0$. The latter case leads to a complication in consumer surplus maximization under decreasing failure rates (see Remark 11). A buyer with type $\theta$ who pays $p$ for a status $s\in[0,1]$ from the good receives a payoff of $u(p,s,\theta)=\theta s-p+v(\theta)$, where $v(\theta)\geq 0$ is the intrinsic value of the good. 888 I assume for now that the intrinsic value is independent of $q$ to abstract from quality differentiation (Mussa and Rosen, 1978). In Section 4.1, I consider an extension where the positional good also has an intrinsic quality equal to ${q}$ produced at a cost $c({q})$. Assume $v^{\prime}(\theta)\geq 0$ and $v^{\prime\prime}(\theta)\leq 0$, so higher types have higher intrinsic values. 999 In Section 4.3, I consider an extension where $v^{\prime}(\theta)\leq-1$, so higher types have lower intrinsic values. If the buyer does not get the good, his status is defined as $s=0$, and he does not get the intrinsic value.
@@
Denote by $G$ the distribution of buyers’ consumption of the good $q$ over $\tilde{X}\equiv X\cup\{0\}$. Then, $G({q})$ represents the mass of buyers who consume weakly lower-level goods than ${q}$ (e.g., smaller cars, lower boarding groups) or opt out ($q=0$). Following Robson (1992) and Hopkins and Kornienko (2004), I define the status of a buyer who consume $q\in X$ as
@@
where $G^{-}({q})\equiv\lim_{{q}\to{q}^{-}}G({q})$ is the mass of buyers who consume strictly lower-level goods than ${q}$ (or opt out). In other words, the status equals the mass of buyers who consume strictly lower-level goods plus half the mass of buyers at the same level.101010 Hopkins and Kornienko (2004) assume a more general specification: $S(q,G(\cdot))=\gamma G({q})+(1-\gamma)G^{-}({q})$ with $\gamma\in[0,1)$. If $\gamma>1/2$ ($\gamma<1/2$), pooling will decrease (increase) the total expected status. I will show that many results are robust to the general specification. An equivalent specification (up to an affine transformation), à la MSS and Dubey and Geanakoplos (2010), is $S({q},G(\cdot))=G^{-}({q})-(1-G(q))\in[-1,1],$ which results in an expected total status of 0. Buyers who opt out ($q=0$) are considered as strictly below those who buy $q\in X$.
@@
Example 2.1 (Priority Services (Gershkov and Winter, 2023)).
A unit mass of customers arrive at the same time. A seller can serve one customer at each instant and offers priority services to skip the line: consumers at higher priority levels are served first, and those of the same level are served in random order. The consumer’s payoff is $u(p,t,\theta)=\tilde{v}(\theta)-\theta t-p$, where $t$ is the waiting time, $\theta$ is the value of time (or the cost of waiting), and $\tilde{v}(\theta)$ is the value of the service. This is equivalent to the model with $s=1-t$ and $v(\theta)=\tilde{v}(\theta)-\theta$. 111111 The assumption $v^{\prime}(\theta)\geq 0$ requires $\tilde{v}^{\prime}(\theta)\geq 1$ here, that is, higher types derive higher utility from the service after accounting for waiting costs. In Section 4.3, I consider an extension where $v^{\prime}(\theta)\leq-1$ (i.e., $\tilde{v}^{\prime}(\theta)\leq 0$) so that higher types derive higher utility from the service after accounting for waiting costs. The seller essentially sells the waiting time $t\in[0,1]$ subject to the feasibility constraint
@@
Because the status category within organization is also a positional good, another example is the status contests in organizations, where the principal (seller) maximizes effort by designing status categories of agents (buyers).
Example 2.2 (Status Contests (Moldovanu et al., 2007)).
Consider a principal who maximizes agent effort by designing status categories. Agents care about their relative positions within the organization. Then, the positional good $q$ represents the status category, the price $p\geq 0$ represents the agent’s effort, and the type $\theta$ represents the agent’s ability (i.e., marginal cost of effort). The agent has a linear effort cost $p/\theta$ and a payoff $s-p/\theta$ if he has status $s\in[0,1]$ in the organization. When the agent is excluded or expelled, he receives zero payoff. This is a special case of the model with $v(\theta)=0$ and $p\geq 0$ after scaling the payoff by $\theta$.
@@ 2.2 A Mechanism Design Approach
Consider a direct mechanism $(p(\theta),\chi(\theta,\omega))$, consisting of a payment function $p\colon\Theta\to\mathbb{R}$ and a (potentially stochastic) positional good allocation $\chi\colon\Theta\times\Omega\to\tilde{X}=X\cup\{0\}$, where the random variable $\omega\in\Omega=[0,1]$ captures the randomness in the mechanism.121212 The allocation can be implemented by a pricing scheme as a function of the distribution of positional good levels. An allocation $\chi$, together with the type distribution $F$, induces a distribution of positional goods $q$, given by $G_{\chi}(q\mid\omega)=\int\mathbf{1}[\chi(\theta,\omega)\leq q]\mathop{}\!\mathrm{d}F(\theta)$.
@@
Denote by $U(\theta)=\theta s-p(\theta)+v(\theta)$ the buyer’s payoff if he participates and $U(\theta)=0$ otherwise. 131313 Although stochastic mechanisms are allowed, it is never optimal to randomly exclude participants either for the seller or consumers. This can be formally shown by multiplying $U(\theta)$ (or $v(\theta)$) by an expected participation dummy $\delta(\theta)\in[0,1]$and showing that $\delta^{*}(\theta)=1$ for all $\theta\geq\theta_{0}$. The following lemmas are implied by standard arguments.
@@
There exists a cutoff type $\theta_{0}\in[0,\bar{\theta}]$ such that the buyer participates (i.e., $U(\theta)\geq 0$) if and only if $\theta\geq\theta_{0}$.
@@
To provide intuitions, consider a deterministic allocation without exclusion $\chi\colon\Theta\to{X}$. Incentive compatibility implies that $\chi(\theta)$ is increasing because $s(\theta)=S(\chi(\theta),G_{\chi})$ is increasing. When $\chi(\theta)$ is strictly increasing at $\theta$, then $s(\theta)=F(\theta)$. Otherwise, when $\chi(\theta)=q>0$ is constant on some interval $[\theta_{1},\theta_{2}]$ (and $\chi(\theta)\neq q$ outside this interval), then
@@
Now I characterize the necessary and sufficient condition for an increasing interim status $s(\theta)$ to be feasible in general.
@@
The proof is à la Theorem 3 (Border’s Theorem) in KMS.
Proof.
[…]
and (iii) fully separating otherwise, where $s(\theta)=F(\theta)$.
Hence, by Proposition 1 (Choquet’s Theorem) in KMS, measure $\lambda_{s}$ supported on the extreme points of $\operatorname{\text{MPS}}_{w}(F)$ such that $s=\operatorname{\mathbf{E}}[\tilde{s}\mid\tilde{s}\sim\lambda_{s}]$. Thus, $s$ can be induced by randomizing over allocations that each induces extreme points of $\operatorname{\text{MPS}}_{w}(F)$. ∎
@@
For deterministic allocations, $s(\theta)$ must be an extreme point of $\operatorname{\text{MPS}}_{w}(F)$—i.e., either the majorization constraint or the monotonicity constraint binds on $[\theta_{0},\theta]$ and $s(\theta)=0$ for all $[0,\theta_{0}]$ (see Theorem 1 and Corollary 2 in KMS). For stochastic allocations, $s(\theta)$ can take other forms.
@@
When exclusion is impossible, the equality must hold at $x=0$, that is, $\operatorname{\mathbf{E}}[s]=\operatorname{\mathbf{E}}[F]=1/2$. In other words, $s(\theta)$ is a mean-preserving spread of $F(\theta)$ in the quantile space, denoted by $s\in\operatorname{\text{MPS}}(F)$.141414 Without exclusion, one can also write $s(\theta)=\operatorname{\mathbf{E}}_{{q}\sim\chi(\theta,\Omega)}[\operatorname{\mathbf{E}}[F(\tilde{\theta})\mid{q}]]$, where $\chi(\theta,\Omega)$ denotes the distribution of $q=\chi(\theta,\omega)$ given $\theta$ generated by the randomness $\omega\in\Omega$.
Remark 3.
Under MSS’s specification $S({q},G)=G^{-}({q})-(1-G({q}))$ without exclusion, the feasibility condition becomes $s\in\operatorname{\text{MPS}}(2F-1)$ with $\operatorname{\mathbf{E}}[s]=\operatorname{\mathbf{E}}[2F-1]=0$.
Remark 4.
In Rayo’s (2013) model of conspicuous goods, where the status value induced by $\chi$ is $\tilde{S}({q},G_{\chi})=\operatorname{\mathbf{E}}[\theta\mid{\chi(\theta)=q}]$, the feasibility condition is $\tilde{s}\in\operatorname{\text{MPS}}_{w}(\theta)$.
Intuitively, we have $s(\theta)=F(\theta)$ under full separation. In contrast, any (partial) pooling of multiple types into the same positional good makes the status $s(\theta)$ a mean-preserving spread of $F(\theta)$ in the quantile space because the value of the positional good becomes lower than its face value, or the value under full separation.
@@
It is straightforward to check that $\operatorname{\mathbf{E}}[s]=1/2$ and that $\int_{\theta}^{\bar{\theta}}s(\theta^{\prime})\mathop{}\!\mathrm{d}F(\theta^{\prime})\leq\int_{\theta}^{\bar{\theta}}F(\theta^{\prime})\mathop{}\!\mathrm{d}F(\theta^{\prime})$ for all $\theta\in[0,\bar{\theta}]$ (with equality at 00, $\theta^{*}$, and $\bar{\theta}$).
@@ 3 Optimal Mechanisms
In this section, I consider the optimal mechanisms that maximize the seller’s revenue and study its welfare implications. Two extreme mechanisms are particularly of interest: full separation (i.e., $s(\theta)=F(\theta)$) and total pooling among participants (i.e., $s(\theta)=(1+F(\theta_{0}))/2$). The former can be induced by offering a continuum of positional goods that lead to a proliferation of status (i.e., the finest partition), and the latter can be induced by offering only one level of positional good.
@@ 3.1 Revenue Maximization
The revenue (or effort) maximization problem is given by
@@
Denote $J(\theta)=\theta-\frac{1-F(\theta)}{f(\theta)}$. Say $F$ satisfies the Myerson’s regularity if $J(\theta)$ is increasing. Say $F$ has an increasing failure rate (IFR) if $\frac{1-F(\theta)}{f(\theta)}$ is decreasing. The revenue is given by
@@
[…]
$\max_{s\in\operatorname{\text{MPS}}_{w}(F),\theta_{0}}\int_{\theta_{0}}^{\bar{\theta}}\left(J(\theta)s(\theta)\right)\mathop{}\!\mathrm{d}F(\theta)+v(\theta_{0})(1-F(\theta_{0})).$ (10)
The maximum is attained at the extreme point of $\operatorname{\text{MPS}}_{w}(F)$, which is an extreme point of $\operatorname{\text{MPS}}(F\cdot\mathbf{1}_{[\theta_{0},\bar{\theta}]})$. The following proposition follows from Fan-Lorentz Theorem (see, e.g., Theorem 4 in KMS).
@@
If $J(\theta)$ is decreasing on some $\Theta_{p}\subseteq[0,\bar{\theta}]$, then the revenue-maximizing mechanism pools every $\theta\in\Theta_{p}$ into the same positional good level.
The proof is in Appendix B.
@@
Proposition 2 still holds under the specification $\tilde{S}(q,G)=\gamma G({q})+(1-\gamma)G^{-}({q})$ with $\gamma\in[0,1/2]$. To see this, when $\gamma<1/2$ (e.g., Immorlica et al. (2015) where $\gamma=0$), pooling is even less likely to be optimal as it reduces expected total status. 161616 Formally, fix a positional good allocation $\chi$, the induced interim status $\tilde{s}(\theta)$ when $\gamma<1/2$ and $s(\theta)$ under the original specification ($\gamma=1/2$) satisfy $\tilde{s}(\theta)\leq s(\theta)$ for all $\theta\in[\theta_{0},\bar{\theta}]$, with equality if and only if $\tilde{s}(x)=F(x)$ for all $x\in[\theta_{0},\theta]$.
Remark 6.
Under the signaling specification $\tilde{S}({q},G_{\chi})=\operatorname{\mathbf{E}}[\theta\mid{\chi(\theta)=q}]$ (see Remark 4), Proposition 2 also holds (with a different optimal cutoff $\theta^{*}_{0}$) because the only difference is that $\tilde{s}\in\operatorname{\text{MPS}}_{w}(\theta)$ instead.
Remark 7.
When restricted to deterministic mechanisms and under the specification $\tilde{S}({q},G)=\phi(S({q},G))$, where $\phi$ is an increasing and convex function, Proposition 2 still holds (with a different $\theta^{*}_{0}$) because Fan-Lorentz Theorem still applies (see details in the proof).
@@
If the virtual valuation $J(\theta)$ is decreasing for some types, the revenue-maximizing mechanism pools these types into the same positional good level. In general, the revenue-maximizing mechanism is derived by applying the ironing technique to obtain the convex hull of $\tilde{J}(\theta)\equiv\int_{0}^{\theta}J(x)\mathop{}\!\mathrm{d}F(x)$.
The proposition has several implications. First, under Myerson’s regularity condition, it is optimal to sell as many levels of positional goods as possible. This is because as the number of position levels increases, the seller can induce a finer partition of buyers and thus extract a higher revenue. 171717 For any two increasing allocations $s,\hat{s}\in\operatorname{\text{MPS}}(F)$ on $[\theta_{0},\bar{\theta}]$, if $s$ is finer than $\hat{s}$, then $\hat{s}\in\operatorname{\text{MPS}}(s)$ (the converse is not true). Thus, I also say $s$ is finer than $\hat{s}$ in terms of majorization if $\hat{s}\in\operatorname{\text{MPS}}(s)$. For example, airlines should introduce more boarding groups (or flight classes) to the extent regulations and practical constraints allow. Luxury companies should offer a proliferation of products from factory stores to high-end.
Second, compared to selling ordinary goods with intrinsic value $v(\theta)$ only, attaching the status value benefits the seller. This explains why luxury companies invest heavily in marketing to create the status value.
Third, the seller also benefits from excluding low types. For uniform distribution on $[0,1]$ and when $v(\theta)=0$, exclusion increases the revenue by 23.5%. Thus, exclusion is important for the seller (organization designer) to extract more revenue (effort). In applications where exclusion is impossible, full separation remains optimal under the regularity condition (See Proposition A.1). The following corollary focuses on the properties of the optimal exclusion level $\theta^{*}_{0}$.
@@
Approximation.
In some applications, the number of positional good levels the seller can offer is subject to regulation, practical constraints, or other costs. The following proposition shows that the seller can still obtain a substantial portion of the maximum revenue by offering only one level of positional good.

### v3

@@ 2.1 Setup
A monopolist seller (she) sells positional goods $q\in X\subseteq\mathbb{R}_{++}$ to a continuum of buyers (he) with unit mass. A larger $q$ represents a higher-level good, such as a larger car or a higher boarding class. Denote by $q=0$ the outside option of not buying the good. Because consumers care about their relative position, consumption of positional goods induces buyers’ status (as defined later). Buyers are heterogeneous in their valuations of status (i.e., types), denoted by $\theta$. Assume the buyer’s type $\theta$ has a distribution $F(\theta)$ with full support $\Theta=[0,\bar{\theta}]$ and density $f(\theta)$. [^8] The support can be extended to $\bar{\theta}=\infty$ and $\underline{\theta}>0$. The latter case leads to a complication in consumer surplus maximization under decreasing failure rates (see Remark 7). A buyer with type $\theta$ who pays $p$ and acquires status $s\in[0,1]$ from the good receives a payoff of $u(p,s,\theta)=\theta s-p+v(\theta)$, where $v(\theta)\geq 0$ is the intrinsic value of the good.  [^9] I assume for now that the intrinsic value is independent of $q$ to abstract from quality differentiation (Mussa and Rosen, 1978). In Section 4.1, I consider an extension where the positional good also has an intrinsic quality equal to q{q} produced at a cost c(q)c({q}). Assume $v^{\prime}(\theta)\geq 0$ and $v^{\prime\prime}(\theta)\leq 0$, so that higher types have higher intrinsic values. If the buyer does not receive the good, his status is defined as $s=0$, and he does not obtain the intrinsic value.
@@
Denote by $G$ the distribution of buyers’ consumption of the good $q$ over $\tilde{X}\equiv X\cup\{0\}$. Then, G(q)G({q}) represents the mass of buyers who consume weakly lower-level goods than q{q} or opt out. Following Robson (1992) and Hopkins and Kornienko (2004), I define the status of a buyer who consume $q\in X$ as
@@
where $G^{-}({q})\equiv\lim_{{q}\to{q}^{-}}G({q})$ is the mass of buyers who consume strictly lower-level goods than q{q} (or opt out). In other words, the status equals the mass of buyers who consume strictly lower-level goods plus half the mass of buyers at the same level. [^10] Hopkins and Kornienko (2004) assume a more general specification: $S(q,G(\cdot))=\gamma G({q})+(1-\gamma)G^{-}({q})$, where $\gamma\in[0,1)$ captures the concern about the number of consumers at the same level. If $\gamma>1/2$ ($\gamma<1/2$), pooling will decrease (increase) the total expected status. I will show that many results are robust to the general specification. An equivalent specification (up to an affine transformation), à la MSS and Dubey and Geanakoplos (2010), is $S({q},G(\cdot))=G^{-}({q})-(1-G(q))\in[-1,1],$ which results in an expected total status of 0. Buyers who opt out ($q=0$) are considered as strictly below those who buy $q\in X$.
@@
Example 1 (Priority Services (Gershkov and Winter, 2023)).
A unit mass of consumers arrive at the same time, and the seller can serve one at each instant. The seller provides priority services to skip the line: consumers at higher priority levels are served first, and those at the same level are served in random order. The consumer’s payoff is $u(p,t,\theta)=\tilde{v}(\theta)-\theta t-p$, where $t$ is the waiting time, $\theta$ is the value of time (or the cost of waiting), and $\tilde{v}(\theta)$ is the value of the service. This is equivalent to the model with s=1ts=1-t and $v(\theta)=\tilde{v}(\theta)-\theta$.  [^11] The assumption $v^{\prime}(\theta)\geq 0$ implies $\tilde{v}^{\prime}(\theta)\geq 1$ here, that is, higher types derive lower utility from the service after accounting for waiting costs. In Section 4.5, I consider an extension where $v^{\prime}(\theta)\leq-1$ (i.e., $\tilde{v}^{\prime}(\theta)\leq 0$) so that higher types derive higher utility from the service after accounting for the waiting cost. The priority services provider essentially sells the waiting time $t\in[0,1]$ subject to the feasibility constraint
@@
Because the status category within organization is also a positional good, another example is the status contests in organizations, where the principal maximizes effort by designing status categories of agents.
Example 2 (Status Contests (Moldovanu, Sela and Shi, 2007)).
Consider a principal who maximizes agents’ effort by designing status categories when they care about their status within the organization. The positional good $q$ represents the status category, the price $p\geq 0$ represents the agent’s effort, and the type $\theta$ represents the agent’s ability (i.e., marginal cost of effort). The agent has a linear effort cost $p/\theta$ and receives a payoff $s-p/\theta$ if he attains status $s\in[0,1]$. When the agent is excluded or expelled, he receives zero payoff. This is a special case of the model with $v(\theta)=0$ and $p\geq 0$ after scaling the payoff by $\theta$.
@@ 2.2 A Mechanism Design Approach
Consider a direct mechanism $(p(\theta),\chi(\theta,\omega))$, consisting of a payment function $p\colon\Theta\to\mathbb{R}$ and a (potentially stochastic) positional good allocation $\chi\colon\Theta\times\Omega\to\tilde{X}=X\cup\{0\}$, where the random variable $\omega\in\Omega=[0,1]$ captures the randomness in the mechanism. [^12] The allocation can be implemented by a pricing scheme as a function of the distribution of positional good levels. An allocation $\chi$, together with the type distribution $F$, induces a distribution of positional goods $q$, given by $G_{\chi}(q\mid\omega)=\int\mathbf{1}[\chi(\theta,\omega)\leq q]\mathop{}\!\mathrm{d}F(\theta)$.
@@
Denote by $U(\theta)=\theta s(\theta)-p(\theta)+v(\theta)$ the buyer’s payoff if he participates and $U(\theta)=0$ otherwise.  [^13] Although stochastic mechanisms are allowed, it is never optimal to randomly exclude participants either for the seller or consumers. This can be formally shown by multiplying $U(\theta)$ (or $v(\theta)$) by an expected participation dummy $\delta(\theta)\in[0,1]$and showing that $\delta^{*}(\theta)=1$ for all $\theta\geq\theta_{0}$. The following lemmas are implied by standard arguments.
@@
There exists a cutoff type $\theta_{0}\in[0,\bar{\theta}]$ such that the buyer participates if and only if $\theta\geq\theta_{0}$.
@@
To provide an intuition, consider a deterministic allocation $\chi\colon\Theta\to\tilde{X}$. Incentive compatibility implies that $\chi(\theta)$ is increasing because $s(\theta)=S(\chi(\theta),G_{\chi})$ is increasing. When $\chi(\theta)$ is strictly increasing at $\theta$, then $s(\theta)=F(\theta)$. Otherwise, when $\chi(\theta)=q>0$ is constant on some interval $\theta_{1},\theta_{2}]$ (and $\chi(\theta)\neq q$ outside this interval), then
@@
Now I provide the necessary and sufficient condition for an increasing interim status $s(\theta)$ to be feasible.
@@
The proof (in Appendix B) is à la Theorem 3 (Border’s Theorem) in KMS.
@@
For deterministic allocations $\chi\colon\Theta\to\tilde{X}$, $s(\theta)$ must be an extreme point of $\operatorname{\text{MPS}}_{w}(F)$—i.e., either the majorization constraint or the monotonicity constraint binds on $\theta_{0},\bar{\theta}]$ and $s(\theta)=0$ for all $0,\theta_{0}]$. For stochastic allocations $\chi$, $s(\theta)$ can take other forms.
@@
When exclusion is impossible, the equality must hold at $x=0$, that is, $\operatorname{\mathbf{E}}[s]=\operatorname{\mathbf{E}}[F]=1/2$. Formally, $s(\theta)$ is a mean-preserving spread of $F(\theta)$ in the quantile space, denoted by $s\in\operatorname{\text{MPS}}(F)$. [^14] Without exclusion, one can also write $s(\theta)=\operatorname{\mathbf{E}}_{{q}\sim\chi(\theta,\Omega)}[\operatorname{\mathbf{E}}[F(\tilde{\theta})\mid{q}]]$, where $\chi(\theta,\Omega)$ denotes the distribution of $q=\chi(\theta,\omega)$ given $\theta$ generated by the randomness $\omega\in\Omega$.  [^15] Under MSS’s specification $S({q},G)=G^{-}({q})-(1-G({q}))$ without exclusion, the feasibility condition becomes $s\in\operatorname{\text{MPS}}(2F-1)$ with $\operatorname{\mathbf{E}}[s]=\operatorname{\mathbf{E}}[2F-1]=0$.
Intuitively, under full separation, the interim status equals the quantile (i.e., $s(\theta)=F(\theta)$). In contrast, (partially) pooling multiple types into the same positional good makes the status $s(\theta)$ a mean-preserving spread of $F(\theta)$ in the quantile space because the value of the positional good becomes lower than its face value—i.e., the value under full separation.
@@
It is straightforward to check that $\operatorname{\mathbf{E}}[s]=1/2$ and that $\int_{\theta}^{\bar{\theta}}s(\theta^{\prime})\mathop{}\!\mathrm{d}F(\theta^{\prime})\leq\int_{\theta}^{\bar{\theta}}F(\theta^{\prime})\mathop{}\!\mathrm{d}F(\theta^{\prime})$ for all $\theta\in[0,\bar{\theta}]$ (with equality at 0, $\theta^{*}$, and $\bar{\theta}$).
@@ 3 Optimal Mechanisms
In this section, I consider the optimal mechanisms that maximize the seller’s revenue and study its welfare implications. Two extreme mechanisms are particularly of interest: full separation (i.e., $s(\theta)=F(\theta)$) and total pooling among participants (i.e., $s(\theta)=(1+F(\theta_{0}))/2$). The former can be induced by offering a continuum of positional goods that lead to a proliferation of status (i.e., the finest partition), and the latter can be induced by offering a single good.
@@ 3.1 Revenue Maximization
The revenue maximization problem is given by
@@
Define $J(\theta)=\theta-\frac{1-F(\theta)}{f(\theta)}$. Say $F$ satisfies the Myerson’s regularity if $J(\theta)$ is increasing. Say $F$ has an increasing failure rate (IFR) if $\frac{1-F(\theta)}{f(\theta)}$ is decreasing. By standard arguments, expected revenue is given by
@@
By (IR), it is optimal to set $U(\theta_{0})=0$. [^16] The price is given by $p(\theta)=\theta_{0}s(\theta_{0})+v(\theta_{0})+\int_{\theta_{0}}^{\theta}x\mathop{}\!\mathrm{d}s(x)\geq 0$. Thus, the revenue maximization problem is equivalent to
$\max_{s\in\operatorname{\text{MPS}}_{w}(F),\theta_{0}}\int_{\theta_{0}}^{\bar{\theta}}J(\theta)s(\theta)\mathop{}\!\mathrm{d}F(\theta)+v(\theta_{0})(1-F(\theta_{0})).$ (10)
The maximum is attained at the extreme point of $\operatorname{\text{MPS}}_{w}(F)$, which is an extreme point of $\operatorname{\text{MPS}}(F\cdot\mathbf{1}_{[\theta_{0},\bar{\theta}]})$. The following proposition follows from the Fan-Lorentz Theorem (see, e.g., Theorem 4 in KMS).
@@
If $J(\theta)$ is not monotonic, define $\tilde{J}(\theta)=\int_{0}^{\theta}J(x)\mathop{}\!\mathrm{d}F(x)$ and let $\operatorname{conv}\tilde{J}$ denote the convex hull of $\tilde{J}$. Then, the revenue-maximizing mechanism pools types into the same positional good level if $\operatorname{conv}\tilde{J}$ is affine and separates types if $\operatorname{conv}\tilde{J}=\tilde{J}$.
@@
Proposition 2 still holds under the more general specification $S(q,G)=\gamma G({q})+(1-\gamma)G^{-}({q})$ with $\gamma\in[0,1/2]$. To see this, when $\gamma<1/2$ (e.g., Immorlica, Stoddard and Syrgkanis (2015) where $\gamma=0$), pooling is even less likely to be optimal as it reduces expected total status.
@@
In general, the revenue-maximizing mechanism is derived by applying the ironing technique to obtain the convex hull of $\tilde{J}(\theta)$. If the virtual valuation $J(\theta)$ is decreasing for some types, the revenue-maximizing mechanism pools them into the same positional good level along with some adjacent types.
The proposition has several implications. First, under Myerson’s regularity condition, it is optimal to sell as many levels of positional goods as possible. This is because, as the number of position levels increases, the seller can induce a finer partition of buyers and thus extract a higher revenue.  [^17] For any two increasing allocations $s,\hat{s}\in\operatorname{\text{MPS}}(F)$ on $\theta_{0},\bar{\theta}]$, if $s$ is finer than $\hat{s}$, then $\hat{s}\in\operatorname{\text{MPS}}(s)$ (the converse is not true). Thus, I also say $s$ is finer than $\hat{s}$ in terms of majorization if $\hat{s}\in\operatorname{\text{MPS}}(s)$. For example, airlines should introduce more boarding groups (or flight classes) to the extent regulations and practical constraints allow. Luxury companies should offer a proliferation of products from factory stores to high-end.
Second, compared to selling ordinary goods with intrinsic value $v(\theta)$ only, attaching the status value benefits the seller. This might explain why luxury companies invest heavily in marketing to create the status value.
Third, the seller also benefits from excluding low types. For uniform distribution on $[0,1]$ and when $v(\theta)=0$, exclusion increases the revenue by 23.5%. Thus, exclusion is important for the seller to extract more revenue. In applications where exclusion is impossible, full separation remains optimal under the regularity condition (See Proposition A.1). The following corollary focuses on the properties of the optimal exclusion level $\theta^{*}_{0}$.
@@
3.1.1 Approximation
In some applications, the number of positional good levels the seller can offer is subject to regulation, practical constraints, or other costs. The following proposition shows that the seller can still obtain a substantial portion of the maximum revenue by offering a single good.